OpenStudy (anonymous):

You deposit $4,000 at the end of each year into an ordinary annuity with an interest rate of 5% compounded annually. What will be the value of the annuity after 10 years? Do not round until the final answer; then round to the nearest dollar as needed.

jimthompson5910 (jim_thompson5910):

Use the formula FV = P((1+r/n)^(n*t) - 1 )/( r/n ) in this case P = 4000 r = 0.05 (5% = 5/100 = 0.05) n = 1 t = 10 plug all this in to get FV = P((1+r/n)^(n*t) - 1 )/( r/n ) FV = 4000((1+0.05/1)^(1*10) - 1 )/( 0.05/1 ) FV = 50311.5701421954 FV = 50311.57

jimthompson5910 (jim_thompson5910):

So the value of the annuity after 10 years is $50,311.57